OpenStudy (anonymous):

show lim (n!)^(1/n) = infinity

OpenStudy (anonymous):

@Hero @ganeshie8 @iambatman

ganeshie8 (ganeshie8):

you may use stirling's approximaiton http://en.wikipedia.org/wiki/Stirling's_approximation

OpenStudy (anonymous):

how about showing lim r^n/n! = 0

ganeshie8 (ganeshie8):

You may use the fact that below series converges to \(e^r\) : \[\sum\limits_{n=0}^{\infty}\dfrac{r^n}{n!}=e^r\] Since the series converges, the corresponding sequence must converge to 0 : \[\lim\limits_{n\to\infty} \dfrac{r^n}{n!}=0\]